f(x)= x^4 +8x^3+16x^2

g(x) = -4x^2 -16x +5

using algebraic method, find the points of intersection of the graphs f(x) and g(x).

which of the solutions are reflected on the graphs


1
Expert's answer
2022-01-19T17:27:04-0500

At point of intersection

"f(x)=g(x)"


"x^4+8x^3+16x^2=-4x^2-16x+5"


"x^4+8x^3+20x^2+16x-5=0"

Factorizing


"(x-0.23606)(x+4.23606)(x^2+4x+5)=0"

Points of intersection are:

x=0.23606

x=-4.23606

This two are for "x\in \reals"

Imaginary points of intersection are:


x={-2-i}

x={-2+i}